# CollabStory: Multi-LLM Collaborative Story Generation and Authorship Analysis

### v2

@@ 5 Authorship Analysis: Extending PAN tasks for multi-LLM scenario
Plagiarism Analysis, Authorship Identification, and Near-Duplicate Detection, known as PAN tasks Bevendorff et al. (2023), have presented a persistent challenge, establishing benchmarks for analyzing multi-authored text among humans for more than 15 years. We extend the most common and repeated authorship-related tasks from the PAN multi-human-author task suite to the multi-LLM scenario. We then fine-tune and report performance using the following 5 baseline methods: Multinomial Naive Bayes (MNB) Losada and Azzopardi (2008), Support Vector Machine (SVM) Vapnik (1998), BERT Kenton and Toutanova (2019), ALBERT Lan et al. (2019), and RoBERTa Liu et al. (2019).
@@ 5.4 Task 4: Authorship Attribution
Authorship Attribution involves predicting exactly who the author of a text article is. In the case of multi-LLM text, we design this task such that each data sample is homogeneous or each part is written by a single author and the classifier’s task is to identify its author. From Figure 3(d), we see that irrespective of the value of $N$, attribution performance is low (between 0.6 to 0.75 F1). We were expecting attribution to be easier the fewer the number of authors in the article since the length of the parts contributed by each author would be longer. But there does not seem to be any such correlation in our dataset i.e. length contributed by each author does not correlate with their detection. We provide more detailed results for each LLM or class-wise F1 scores in the Appendix (Table 8).
@@ A.1 LLM Prompting
For for all $N\geq 2$, we provided the summary of the story so far as an input in the prompt. To make sure that our story parts had smooth continuity, we also used the last sentences of the story so far as input. This made sure that the generating LLM has access to the last sentence in addition to the overall storyline to continue the story as seamlessly as possible. This second input is denoted as “last sentence from previous part” in Table 3. Other than this, our prompt for the three types of story sections only differed in the instruction of writing either the “beginning”, “continue”, or “conclude” the storyline so far. We also had to add an instruction to stop the LLMs from generating any additional instructions as from our pilot study, we found that some LLMs (Orca and Llama) would often first generate a rephrasing or more detailed version of our instruction before generating the actual story content.
@@ Appendix B Authorship Analysis: Real world implications
This work have profound implications for various stakeholders in the burgeoning socio-technical system of generative AI. Our research introduces authorship-related tasks using CollabStory, which can help address these concerns by accurately discerning the usage of multiple LLMs in texts (summarized in Table 6. Our extension of PAN-inspired authorship tasks is closely linked to real-world implications, as follows:

### v3

@@ 5 Authorship Analysis: Extending PAN tasks for multi-LLM scenario
Plagiarism Analysis, Authorship Identification, and Near-Duplicate Detection, known as PAN tasks Bevendorff et al. (2023), have presented a persistent challenge, establishing benchmarks for analyzing multi-authored text among humans for more than 15 years. We extend the most common and repeated authorship-related tasks from the PAN multi-human-author task suite to the multi-LLM scenario. We then fine-tune and report performance using the following 5 baseline methods: Multinomial Naive Bayes (MNB) Losada and Azzopardi (2008), Support Vector Machine (SVM) Vapnik (1998), BERT Kenton and Toutanova (2019), ALBERT Lan et al. (2019), and RoBERTa Liu et al. (2019). We used a 70:30 split of the dataset for fine-tuning and evaluation and aggregated performance over three random splits.
@@ 5.4 Task 4: Authorship Attribution
Authorship Attribution involves predicting exactly who the author of a text article is. In the case of multi-LLM text, we design this task such that each data sample is written by a single author (does not contain any text windows that have an authorship change), and the classifier’s task is to identify its author. From Figure 3(d), we see that irrespective of the value of $N$, attribution performance is low (between 0.6 to 0.75 F1). We were expecting attribution to be easier the fewer the number of authors in the article since the length of the parts contributed by each author would be longer. But there does not seem to be any such correlation in our dataset i.e. length contributed by each author does not correlate with their detection. We provide more detailed results for each LLM or class-wise F1 scores in the Appendix (Table 9).
@@ 6 Conclusion
7 Acknowledgments
This work was in part supported by NSF awards #1934782 and #2114824. Some of the research results were obtained using computational resources provided by CloudBank (https://www.cloudbank.org/), which was supported by NSF award #1925001.
@@ A.1 LLM Prompting
For all $N\geq 2$, we provided the summary of the story so far as an input in the prompt. To make sure that our story parts had smooth continuity, we also used the last sentences of the story so far as input. This ensured that the generating LLM has access to the last sentence and the overall storyline to continue the story as seamlessly as possible. This second input is denoted as “last sentence from previous part” in Table 3. Other than this, our prompt for the three types of story sections only differed in the instruction of writing either the “beginning”, “continue”, or “conclude” the storyline so far. We also had to add an instruction to stop the LLMs from generating any additional instructions as from our pilot study, we found that some LLMs (Orca and Llama) would often first generate a rephrasing or more detailed version of our instruction before generating the actual story content.
@@ Appendix B Authorship Analysis: Real world implications
This work has profound implications for various stakeholders in the burgeoning socio-technical system of generative AI Lau et al. (2024); van der Sloot (2024); Franceschelli et al. (2024); Lucas et al. (2024). Our research introduces authorship-related tasks using CollabStory, which can help address these concerns by accurately discerning the usage of multiple LLMs in texts (summarized in Table 7. Our extension of PAN-inspired authorship tasks is closely linked to real-world implications, as follows: